# Polarized multiwavelength emission from pulsar wind – accretion disk interaction in a transitional millisecond pulsar

### v1

@@ A.5 VLA
We created Stokes images with the WSClean imager [94]; WSClean is a CLEAN-based deconvolution algorithm [66] that removes sampling artefacts from interferometric images. To produce our final data products, we first performed a shallow unmasked deconvolution of the full 10′ field of view (FOV), i.e., we created preliminary images. We made a deconvolution mask using the unmasked Stokes image and, with the mask, re-imaged the target field. Adopting the masked image as a model, we performed phase-only self-calibration on the target visibilities with the software package QuartiCal [75]. We then refined the mask and repeated the imaging and self-calibration procedure a second time. In the FOV of J1023, there exists, among other fainter background sources, a bright radio Galaxy (J102358.2003826) at an angular offset of . Using the self-calibrated model, we subtract out the visibilities for the background sources, allowing for narrow FOV images that are minimally affected by aliasing.
During the VLA observations, J1023 had an average flux density of 240 Jy at 6 GHz. We created two sets of images using our background-subtracted visibilities. First, we broke the observations into 1-minute imaging intervals, allowing us to (temporally) resolve the radio variability. The 1-minute images revealed that J1023 underwent a mini-flare lasting 8 min and reaching a maximum flux density of 500 Jy. We created four additional images that combined multiple target scans to get deeper polarization limits, sacrificing temporal resolution for higher signal-to-noise. The deep imaging time intervals were 23:10:31–23:52:25 UTC (June 4), 23:52:25–00:00:22 UTC (June 4 to June 5) and 00:00:22.5–00:36:20 UTC (June 5); these intervals correspond to the times before the mini-flare, the mini-flare, and the times after the mini-flare. The fourth image combined the times before and after the mini-flare.
In each imaging interval, WSClean outputted 16 channelized images (evenly spaced in frequency) and a single frequency-averaged ‘Multi-Frequency-Synthesis’ (MFS) image. We measure the target’s Stokes flux density from the MFS image, using the CASA task imfit. imfit models the intensity distributions as an elliptical Gaussian; given that J1023 is a point source, we enforce a synthesized beam-shaped Gaussian fit. We quantified the uncertainty on the flux density with the root-mean-square (RMS) noise extracted from a nearby source-free region. Using the CASA task imstat, we measure the RMS using a circular aperture with an area equal to 100 synthesized beam area. Additionally, for MFS images where J1023 was detected at (corresponding to in each channelized image), we calculate the intra-band power-law spectral index from channelized flux densities (using scipy.curvefit) , which describes how the flux density varies with observing frequency as . We found that remains nearly constant during the high-mode episodes that precede (H1) and follow (H2) the low mode, but decreases during the low mode (L). Specifically, we obtain , and .
Lastly, we produced linear polarization intensity images for each imaging interval but did not detect any polarized emission. As a result, we calculate 3 ( 99.7%) percent upper-limits from the -images following the prescription by [120]. We obtain , and . By stacking high-mode data, we obtain .
Appendix B Selection of Emission Modes in the IXPE Dataset
@@
To select the appropriate thresholds for the count rates in order to select the distinct emission modes in the IXPE data, we proceeded as follows. Over the past decade, the net count rate observed in all previous observations of J1023 carried out by the European Photon Imaging Cameras (EPIC) on board XMM-Newton decreased by a factor of approximately six during the low mode compared to the average net count rate (e.g., [18, 33]). The average count rate of J1023 measured by IXPE after combining data from the three DUs is 0.09 counts s-1, with about 30% of this rate contributed by the background level. Consequently, during low-mode episodes, typically only 0 or 1 net photon from the source, plus 3–4 photons from the background, would be detected by IXPE in 100-s time bins. In fact, during the low modes covered by NICER, the highest count rate measured in the IXPE time series binned at 100 s is 0.05 counts s-1 (summing up the source and background contributions). We then classified time intervals in the IXPE light curves as low modes whenever the count rate in the 100-s binned time series was less than or equal to 0.05 counts s-1. This criterion resulted in the selection of about 182 ks of low-mode data, which corresponds to about 27% of the total exposure time. This value is consistent with the fraction of time that J1023 was detected in the low mode in previous observations [5, 18, 33].
@@
To differentiate between high and flaring modes, we extracted the distribution of count rates above 0.05 counts s-1 and modeled it using two Gaussian functions. We then classified a 100-s time bin as a flare mode if its count rate was at least 4 standard deviations above the mean count rate of the Gaussian distribution representing the high mode, which corresponds to a threshold of 0.2 counts s-1 (Fig. 10). This approach resulted in the selection of 20 ks of flaring-mode data, which represents 3% of the total duration of the IXPE observations. This percentage is again consistent with the fraction observed in previous observations [5, 18, 33].
Appendix C Timing Analysis of the IXPE Dataset
Previous studies have shown that coherent X-ray pulsations at the NS spin period in J1023 are detected only during the high mode [5, 71, 98, 70]. Hence, to search for a pulsed signal in the IXPE data, we retained only source photons collected during the high mode, and restricted the analysis to photons in the energy range 2–6 keV. We corrected the photons’ arrival times for the motion of the pulsar along the binary orbit, setting the orbital period and the projected semi-major axis of the circular orbit to the values reported by [3], = 17115.5216592 s and asini/c = 0.343356 light-seconds (the uncertainties in these parameters are of the order 10-7 s and 10-6 light-seconds, respectively). To measure the epoch of passage of the pulsar at the ascending node of the orbit (), we performed an epoch folding search on the IXPE dataset around the spin frequency value expected according to the previously measured secular evolution [71, 22], sampling each spin cycle in 16 phase bins. We used a grid of values centered on the value predicted according to the long-term evolution measured by [70], MJD, with a spacing of 0.2 s. We recovered the coherent signal with an epoch folding variance of (corresponding to a single-trial false alarm rate of ) at a spin frequency of Hz and = 60461.979580(14) MJD. The uncertainty on the spin frequency was evaluated following [80], while that on was taken as the half-width at half-maximum of the Gaussian fit to the pulse variance distribution. The timing model, compatible with the Tempo2 [65] and PINT [86] pulsar timing packages, will be made available at Zenodo.

### v2

@@ 2.5 VLA Observations
Data were processed using the CASA VLA pipeline (CASA Team et al., 2022a), followed by manual and automated (rflag and tfcrop) radio frequency interference flagging. We performed cross-hand calibration following standard VLA polarization procedures444https://casaguides.nrao.edu/index.php/CASA_Guides:Polarization_Calibration_based_on_CASA_pipeline_standard_reduction:_The_radio_galaxy_3C75-CASA6.5.4. We created Stokes images with the WSClean imager (A. R. Offringa et al., 2014); WSClean is a CLEAN-based deconvolution algorithm (J. A. Högbom, 1974) that removes sampling artefacts from interferometric images. To produce our final data products, we first performed a shallow unmasked deconvolution of the full 10′ field of view (FOV), i.e., we created preliminary images. We made a deconvolution mask using the unmasked Stokes image and, with the mask, re-imaged the target field. Adopting the masked image as a model, we performed phase-only self-calibration on the target visibilities with the software package QuartiCal (J. S. Kenyon et al., 2022). We then refined the mask and repeated the imaging and self-calibration procedure a second time. In the FOV of J1023, there exists, among other fainter background sources, a bright radio Galaxy (J102358.2003826) at an angular offset of . Using the self-calibrated model, we subtract out the visibilities for the background sources, allowing for narrow FOV images that are minimally affected by aliasing.
We first produced time-resolved images in 1-min intervals. These images revealed that J1023 experienced a short (8 min) mini-flare peaking at 500 Jy. We then created also four deeper polarization images by combining multiple scans to boost signal-to-noise ratio at intervals corresponding to the times before the mini-flare, the mini-flare, the times after the mini-flare, and the combined times before and after the mini-flare. For each interval, WSClean generated 16 channelized images and a single frequency-averaged Multi-Frequency Synthesis (MFS) image. Flux densities were measured using imfit in CASA, modeling the source as a point source with a synthesized beam-shaped Gaussian fit. uncertainties were derived from RMS noise estimates in nearby source-free regions using imstat, employing a circular aperture encompassing 100 synthesized beam areas. For MFS images where J1023 was detected at (corresponding to in each channelized image), we fitted a power-law of the form to channelized flux densities to determine the intra-band spectral index .
Lastly, we produced linear polarization intensity images for each imaging interval.
3 Data Analysis and Results
3.1 Timing Analysis of the IXPE Dataset
Coherent X-ray pulsations at the NS spin period in J1023 are detected only during the high mode (A. M. Archibald et al., 2015; A. Jaodand et al., 2016; A. Papitto et al., 2019; G. Illiano et al., 2023). Hence, we used source photons collected by IXPE in the high mode in the 2–6 keV band for our analysis. We corrected photon arrival times for the pulsar’s orbital motion, using an orbital period of = 0.1980963155 d and a projected semi-major axis of asini/c = 0.343356 light-seconds (G. Illiano et al., 2023). To measure the epoch of passage of the pulsar at the ascending node of the orbit (), we performed an epoch folding search around the spin frequency value expected according to the previously measured secular evolution (A. Jaodand et al., 2016; A. Burtovoi et al., 2020), sampling each spin cycle in 16 phase bins. We used a grid of values centered on the value predicted according to the long-term evolution measured by G. Illiano et al. (2023), MJD, with a spacing of 0.2 s. We recovered the coherent signal with an epoch folding variance of (single-trial false alarm rate ) at a spin frequency Hz and = 60461.979580(14) MJD. The uncertainty on the spin frequency was evaluated following D. A. Leahy (1987), while that on was taken as the half-width at half-maximum of the Gaussian fit to the pulse variance distribution555The timing model is archived at Zenodo.. We obtained compatible results by analysing 0.5–6 keV NICER data in the high mode taken on 2024 June 1. The timing parameters had slightly larger uncertainties due to the lower number of photons recorded () compared to IXPE () and the uneven coverage. Our measurements tend to favor a scenario where the pulsar’s rotational evolution in the subluminous X-ray state remains largely unchanged compared to its behavior during the radio pulsar state (Fig. 8).
3.2 X-ray Polarimetric Properties
Details on the polarimetric analysis of the IXPE data are given in Appendix C. Within the 3–6 keV band, we detected a polarization degree of % and a polarization angle , measured counter-clockwise from the North celestial pole toward the East. Hereafter, all uncertainties are reported at 68% c.l.. This polarization degree exceeds the Minimum Detectable Polarization (MDP; M. C. Weisskopf et al. 2010; R. F. Elsner et al. 2012) of 15% at 99% confidence, allowing us to reject the null hypothesis of unpolarized emission with 99.4% confidence (3.2 significance666https://heasarc.gsfc.nasa.gov/docs/ixpe/analysis/IXPE_Stats-Advice.pdf). In the 2–3 keV band, the polarization degree is consistent with zero within a 1 uncertainty. A similar analysis of the high-mode data yields values consistent with those reported above, though none exceeded the MDP at 99% confidence. No significant variations in Stokes parameters were observed across the pulsar rotational cycle or the binary orbital cycle (Appendix C.3). For low-mode and flaring-mode data, the significance was too low for a reliable analysis, yielding upper limits of % and % at 90% confidence over the 2–6 keV range (Appendix C.1).
To improve the polarimetric measurements, we performed weighted spectro-polarimetric analyses (T. E. Strohmayer, 2017; A. Di Marco et al., 2022) on both the entire dataset and the high-mode dataset (Appendix C.2). We fitted a power-law model corrected for interstellar absorption (S. Bogdanov et al., 2015; F. Coti Zelati et al., 2018) and convolved with a constant polarization model to the data in the 2–6, 2–3 and 3–6 keV ranges. Results are summarized in Table 2. For the average emission in the 2–6 keV range, we obtained % at . For the high-mode emission in the same range, we measured % at . To estimate the significance level, we used the steppar command in xspec on PD and PA. The contour remains open at a significance above 99.7%. Figure 2 shows the confidence contours for the measurements of and in the high mode across energy bands.
Previous analyses show that around 10% of low-mode episodes in the IXPE datasets may be missed due to their short durations (A. M. Archibald et al., 2015; S. Bogdanov et al., 2015; F. Coti Zelati et al., 2018). However, this does not affect the polarization degree measured during the high mode, as the potential increase, about 0.2% in the polarization degree, is within the measurement uncertainties (see Appendix C.2).
| Energy Band (keV) | Average Emission | High Mode | ||||
| (%) | (∘) | /dof | (%) | (∘) | /dof | |
| 2–6 | 175/166 | 164/166 | ||||
| 2–3 | 32/31 | 25/31 | ||||
| 3–6 | 120/121 | 120/121 | ||||
3.3 Optical Properties
Details on the optical properties of J1023 are presented in Appendix D. J1023 was detected at an average magnitude of 16.4 (Vega) in the -band. We measured an average linear polarization degree of and a polarization angle of . During the X-ray high mode, we measured at an angle . This polarization angle is fully consistent with that derived at X-ray energies within the uncertainties (Fig. 2). During the X-ray low mode, the polarization degree decreases slightly to . However, the polarization angles in high and low modes remain statistically consistent with each other.
3.4 Radio Properties
J1023 exhibited an average flux density of 240 Jy at 6 GHz and experienced a mini-flare matching a low-mode episode in X-rays lasting 8 min, reaching a peak flux density of 500 Jy (Fig. 3). The power-law spectral index remains nearly constant during the high-mode episodes that precede (H1) and follow (H2) the low mode, but decreases during the low mode (L): we obtain , and . These results are consistent with previous observations (S. Bogdanov et al., 2018) which pointed to optically thick synchrotron emission in the high mode and a mix of optically thick and optically thin synchrotron emission in the low mode (M. C. Baglio et al., 2023).
No radio polarization was detected (Fig. 3). Therefore, we derived 3 ( 99.7%) upper limits on the polarization fraction from the -images following the prescription outlined by J. E. Vaillancourt (2006). We established , and . By stacking high-mode data, we achieved . These results represent the first constraints on the linear polarization fraction of the radio emission in both modes.
3.5 Spectral Energy Distributions
We extracted the SEDs for the high mode (see Fig. 4) by combining the unabsorbed X‑ray fluxes in the 2–3, 3–4 and 4–6 keV energy bands from IXPE with the averaged -band fluxes from VLT/FORS2. Adopting the estimated hydrogen column density (), we calculated the -band absorption coefficient based on the relation from D. R. Foight et al. (2016) and determined the -band absorption using the extinction laws from K. D. Gordon et al. (2023) (see also K. D. Gordon et al. 2009; E. L. Fitzpatrick et al. 2019; C. Gordon & K. Arnaud 2021; M. Decleir et al. 2022).
For the polarized SED, we used the polarization degrees measured in the 2–3 keV and 3–6 keV bands through spectro-polarimetric analyses (see Table 2). The dereddened polarized optical flux was estimated by rescaling according to the polarization degree in the high mode, .
For comparison, we overlaid the total and pulsed X-ray and optical fluxes obtained from a previous campaign involving the XMM-Newton and NuSTAR satellites as well as the SiFAP2 photometer mounted on the INAF Telescopio Nazionale Galileo (A. Papitto et al., 2019). Additionally, we included the fit of their SED of the pulsed emission using a single power law defined by , where is the flux density at frequency , is the normalization constant, and is the spectral index. Using nonlinear least squares fitting, the best-fit parameters were determined to be , and (A. Papitto et al., 2019).
Our analysis shows that the SED of the polarized flux is consistent with the same power-law model that describes the spectrum of the pulsed signal. This provides additional strong evidence that the observed polarized emission shares a common origin with the pulsed emission across both optical and X-ray bands. Moreover, the near-coincidence of the polarized and pulsed fluxes throughout the spectrum indicates that the observed polarization originates almost entirely from the same emission processes responsible for the pulsations across the optical and X-ray frequencies. Theoretical predictions for the linear polarization of synchrotron radiation, assuming a spectral index of (as measured by A. Papitto et al. 2019 and shown in Fig. 4), suggest a theoretical maximum polarization degree of 77% (G. B. Rybicki & A. P. Lightman, 1979), which is consistent with our findings. Future instrumentation designed to measure the polarization of pulsed emission will be essential to test this.
4 Discussion
4.1 The Nature of the Polarized Emission
4.1.1 An Accretion-powered Mechanism?
Theoretical models for accreting NSs in LMXBs predict that the primary X-ray emission mechanisms (including thermal Comptonization, emission from the spreading/boundary layer, disk reflection, and fan-beamed radiation from small, heated surface spots) should yield polarization degrees of only a few percent (for a review, see F. Ursini et al. 2024). For example, standard Comptonization models predict that seed photons scattering off high‑energy electrons yield only modest polarization. Similarly, theoretical studies constrain the polarization from emission in the spreading layer to be 1.5% (A. Bobrikova et al., 2023, 2024; R. Farinelli et al., 2024). These theoretical expectations are well supported by the IXPE observations conducted so far on high-luminosity accreting NSs (with X-ray luminosities at least three orders of magnitude higher than J1023), where these processes resulted in polarization degrees of only a few percent (see A. Papitto et al. 2025 for the case of an accreting MSP).
In stark contrast, our observations of J1023 show a substantially higher polarization degree in the X-ray band. Furthermore, the pure power-law spectrum of J1023 lacks additional components (e.g., multiple Comptonized components, disk reflection features, or signatures of localized fan-beamed emission) that would be expected if the above-mentioned standard mechanisms were significantly contributing to the X-ray emission. This discrepancy strongly suggests that an additional or alternative mechanism must be at work to generate the much higher polarization degree observed from J1023.
Scattering of emission by outflows of highly ionized material launched from the accretion disk along the equatorial plane – commonly referred to as disk winds (e.g. M. Díaz Trigo & L. Boirin 2016) – may also provide a source of polarization. Although disk winds have not been observed in J1023, recent work (A. P. Nitindala et al., 2025) suggests that even if they were present and contributed significantly to the X‑ray emission, such winds alone could not account for the high polarization degree given the system’s orbital inclination of ; J. G. Stringer et al. 2021).
4.1.2 A Rotation-powered Mechanism?
A few models have been developed to explain the X-ray to optical polarization properties in isolated pulsars, including the outer gap and two-pole caustic models, as well as current sheet models (for a review, see A. K. Harding 2019). In what follows, we focus on our IXPE results in the 2–6 keV band (see Fig. 11) and compare them with the predictions from these models, keeping in mind that the photon statistics in our dataset limit the degree to which we can definitively confirm or rule out any particular scenario.
The outer gap (OG; K. S. Cheng et al. 1986, 2000; A. K. Harding & A. G. Muslimov 2001; J. Takata et al. 2008) and two-pole caustic (TPC; J. Dyks & B. Rudak 2003) models propose that the high-energy emission of pulsars originates from the outer magnetosphere, along the last open magnetic field lines inside the light cylinder radius. The emission mechanisms involve synchrotron radiation and curvature radiation from particles accelerated within vacuum gaps or extended regions of strong electric fields. Pulse profiles are characterized by caustics formed due to relativistic effects like aberration and light travel time delays. This results in double-peaked light curves, with peak separation depending on the magnetic inclination angle and the observer’s viewing angle. These models predict polarization properties such as rapid swings in the PA correlated with the pulses and dips in PD at the pulse peaks. The dips are attributed to depolarization from overlapping emissions originating from regions with varying magnetic field orientations. In the TPC model, rapid PA swings are expected through both peaks of double-peaked pulse profiles. In contrast, the OG model predicts a significant PA swing primarily at the second peak, and the first peak may not show a strong PA swing unless the emission occurs very near the light cylinder (J. Dyks et al., 2004).
The X-ray PA in J1023 remains relatively constant across the pulse profile within the uncertainties, showing variations by a few tens of degrees at most (Fig. 11). This is at odds with the strong, rapid PA swings of up to 180∘ predicted by OG/TPC models (A. K. Harding, 2019).
Current sheet models (J. Pétri & J. G. Kirk, 2005; B. Cerutti et al., 2016) focus on emission from particles accelerated in the equatorial current sheet beyond the light cylinder. Magnetic reconnection in the current sheet accelerates particles, leading to high-energy synchrotron emission. Pulses are formed when the observer’s line of sight crosses the current sheet, resulting in one or two pulses per rotation. These models predict PD values typically 15–30%, with PD dips at the pulse peaks due to depolarization effects. Significant PA swings are expected at the pulse peaks due to the reversal of the magnetic field polarity across the current sheet.
J1023 does not exhibit the sharp X-ray PA swings predicted by current sheet models (Fig. 11). This lack of pronounced PA variation suggests a more uniform magnetic-field orientation in the emission region, inconsistent with the magnetic field reversals expected in current sheet models.
Overall, we find no clear evidence for the pronounced polarimetric swings predicted by standard rotation-powered pulsar emission models. Although the limited photon statistics in some phase bins prevent us from excluding milder geometric variants of these models, the absence of prominent polarimetric swings at pulse maxima in the X-ray band is more consistent with a relatively uniform magnetic-field configuration rather than with the dramatic field reversals typically required by such models.
4.1.3 Pulsar Wind – Accretion Flow Interaction?
A scenario for the high mode of J1023 suggests the presence of an active rotation-powered pulsar, an accretion disk, and a compact jet. In this scenario, the observed X-ray, UV, and optical pulses originate from synchrotron emission at a boundary region where the pulsar electromagnetic wind – modulated at the NS spin period – collides with the inner accretion flow (A. Papitto et al., 2019; A. Veledina et al., 2019). This interaction is thought to occur just beyond the light cylinder radius (the distance at which corotation with the pulsar becomes physically impossible), which is 80 km for J1023. The observed lag of 150 s between optical and X-ray pulses is suggested to originate from differences in the cooling timescales of synchrotron photons at these energies, given the expected magnetic field strength at the boundary region (A. Papitto et al., 2019; G. Illiano et al., 2023). A compact jet is launched likely near the compact object, emitting partially self-absorbed synchrotron radiation from radio to infrared frequencies. Previous work shows that the emission at the boundary region contributes 3% of the optical flux and 83% of the 0.3–10 keV flux, while the compact jet emission accounts for 2.5 in the optical band and 17% in the 0.3–10 keV energy band (M. C. Baglio et al., 2023).
Electrons spiraling along toroidal magnetic field lines around the jet axis emit synchrotron radiation, with the polarization oriented perpendicular to the field lines. Given this magnetic field geometry and the orbital inclination of J1023, calculations indicate that the expected polarization degree should not exceed 25% (A. Veledina & M. Pélissier, 2025; M. Lyutikov et al., 2005; H. Krawczynski et al., 2022). In the -band, this translates to a maximum polarization degree of 0.6%, which is lower than the average optical polarization observed for J1023. In the IXPE energy band, the limit is 3%, again below our measured value. Hence, based on these considerations, we rule out the compact jet as the main source of the observed polarization.
If the observed low-level optical polarization is due to synchrotron radiation from the boundary region, and a single population of electrons accelerated at this region produces synchrotron radiation from optical to X-rays, then the intrinsic polarization in the optical band should extend to higher energies. In this scenario, we estimate an intrinsic polarization degree of 12–17% in the X-ray band and expect close alignment between the average optical and X‑ray polarization angles (M. C. Baglio et al., 2023). Our measurements are fully consistent with these predictions, providing striking evidence that the polarized emission in J1023 is driven by emission at the boundary region.
Figure 4 shows the SEDs of the high-mode flux, polarized flux, and pulsed flux of J1023 across optical and X-ray wavelengths. The SED of the polarized flux follows the same power law that describes the SED of the pulsed flux. Moreover, both the polarized and pulsed fluxes are consistent with each other. These results suggest that the polarized and pulsed emissions share a common origin, lending support to the scenario in which the synchrotron-emitting boundary region is the primary source of both in J1023.
Figure 5 illustrates the geometry of the system in the high mode. The of synchrotron radiation traces the orientation of the magnetic field projected onto the sky plane, perpendicular to the observer’s line of sight. If the observed pulsed emission from optical to X-ray energies arises from synchrotron radiation emitted by particles accelerated at regions on the NS wind – disk boundary that corotate with the NS (A. Papitto et al., 2019), the magnetic field geometry strongly influences how the varies with pulsar rotation. A predominantly toroidal magnetic field at the emission site would produce a periodic modulation of the by 180∘ each rotation. Conversely, a predominantly poloidal magnetic field configuration would yield a nearly constant, phase-independent (A. Veledina & M. Pélissier, 2025). Although a modest toroidal component cannot be entirely ruled out, given statistically significant variations observed in a few pulse-phase bins, the overall consistency of the across the spin cycle (Fig. 11) suggests a predominantly poloidal field geometry. Because the observed synchrotron emission integrates contributions from many small regions across the wind–disk boundary – each characterized by similar magnetic-field orientations and electron energy distributions – local fluctuations average out, producing a that remains nearly constant throughout the pulsar rotation.
4.2 Synchrotron Emission across Scales
Our findings reveal striking similarities between J1023 and pulsar wind nebulae (PWNe). In PWNe, the high-energy emission originates from synchrotron radiation by electrons accelerated within highly ordered magnetic fields. These fields are shaped by the pulsar wind as it wraps around the pulsar and is compressed within the nebula (see, e.g., S. P. Reynolds et al. 2012). However, PDs in PWNe typically exceed those measured in J1023. For example, the Vela PWN exhibits a X-ray PD averaging 45%, reaching up to 70% in inner regions (F. Xie et al., 2022). Additionally, a PD exceeding 60% has been detected in the outskirts (K. Liu et al., 2023). The Crab PWN shows a spatially integrated PD of 20% (N. Bucciantini et al., 2023), with some regions exhibiting PDs of 45 – 50% (K. Liu et al., 2023; N. Bucciantini et al., 2023; C. Zuo et al., 2025). The PWN around the pulsar PSR B054069 reaches a PD of 25% (F. Xie et al., 2024). The PWN 3C 58 is characterized by an integrated PD of 22% (N. Bucciantini et al., 2025).
The distinct geometries and spatial scales of PWNe compared to J1023 may play a crucial role in the observed differences. In PWNe, the pulsar wind expands into the surrounding environment over large distances, with the wind termination shock occurring at 0.1 pc from the pulsar (J. J. Hester et al., 2002; R. Petre et al., 2007). This extended region facilitates the development of large-scale, toroidal magnetic fields with highly ordered structures. At the same time, it disperses any coherent pulsations emitted by the NS. Consequently, PWNe are not expected to produce pulsed emission through the same emission mechanism described for J1023. In contrast, J1023 features a much more compact emission region located 100 km from the pulsar. At such close proximity, the interaction between the pulsar wind and the inflowing material from the accretion disk is more disruptive to the striped wind structure due to the higher density and turbulence of the accretion flow. This interaction likely results in a more tangled magnetic field at the boundary region, which may be the main factor accounting for the lower polarization degree observed in J1023 compared to PWNe.
An especially relevant comparison comes from recent IXPE observations of the binary pulsar PSR B125963 (P. Kaaret et al., 2024), where the pulsar wind interacts with a massive companion’s decretion disk. Despite the different source of the interacting matter, PSR B125963 shows a PD (8%) similar to J1023. This similarity reflects a common scenario in binary systems, where the interactions between pulsar winds and dense matter flows in compact and turbulent boundary regions inherently produce less ordered magnetic fields and thus reduced polarization in the emitted radiation, compared to the highly structured fields seen in PWNe.
5 Conclusions
The SEDs of the polarized and pulsed fluxes align with a single power-law model across optical and X-ray energies. This suggests that the polarized emission originates from the same processes responsible for the pulsed emission.
Our findings support a scenario where the polarized emission is produced by synchrotron radiation at a boundary region formed by the interaction between the pulsar wind and the inner accretion disk. This gives insights into how the pulsar wind interaction with its environment shapes the observed emission properties.
These findings enhance our understanding of the complex interplay between accretion-powered and rotation-powered processes in tMSPs, and establish multiwavelength polarimetry as a crucial diagnostic tool for uncovering the physical mechanisms at work in these systems. Future observations, particularly those capable of measuring the polarization of the multiwavelength pulsed emission, will be essential to further elucidate the nature of the emission processes in J1023 and similar objects.
Appendix A Selection of Emission Modes in the IXPE Dataset
@@
To select the appropriate thresholds for the count rates in order to select the distinct emission modes in the IXPE data, we proceeded as follows. Over the past decade, the net count rate observed in all previous observations of J1023 carried out by the European Photon Imaging Cameras (EPIC) on board XMM-Newton decreased by a factor of approximately six during the low mode compared to the average net count rate (e.g., S. Bogdanov et al. 2015; F. Coti Zelati et al. 2018). The average count rate of J1023 measured by IXPE after combining data from the three DUs is 0.09 counts s-1, with about 30% of this rate contributed by the background level. Consequently, during low-mode episodes, typically only 0 or 1 net photon from the source, plus 3–4 photons from the background, would be detected by IXPE in 100-s time bins. In fact, during the low modes covered by NICER, the highest count rate measured in the IXPE time series binned at 100 s is 0.05 counts s-1 (summing up the source and background contributions). We then classified time intervals in the IXPE light curves as low modes whenever the count rate in the 100-s binned time series was less than or equal to 0.05 counts s-1. This criterion resulted in the selection of about 182 ks of low-mode data, which corresponds to about 27% of the total exposure time. This value is consistent with the fraction of time that J1023 was detected in the low mode in previous observations (A. M. Archibald et al., 2015; S. Bogdanov et al., 2015; F. Coti Zelati et al., 2018).
@@
To differentiate between high and flaring modes, we extracted the distribution of count rates and modeled it using Gaussian mixture models. We then classified a 100-s time bin as a flare mode if its count rate was at least 4 standard deviations above the mean count rate of the Gaussian distribution representing the high mode, which corresponds to a threshold of 0.2 counts s-1 (Fig. 7). This approach resulted in the selection of 20 ks of flaring-mode data, which represents 3% of the total duration of the IXPE observations. This is again consistent with the fraction observed in previous observations (A. M. Archibald et al., 2015; S. Bogdanov et al., 2015; F. Coti Zelati et al., 2018).
Appendix B Timing Analysis of the IXPE Dataset